hello I have got a question from my homework that states that i am a lab technician that needs to make 60ml of 32% nitric acid solution. The acid solutions available are 20% nitric acid solution and 40% nitric acid solution.How many milliliters of each solution should be mixed to make the 32% solution. I have read this question several times and don't know what they are asking me to do and i cannot understand how to get the solution. I hope you understand my question i have tried to make it very clear and am sorry if this is considered -dump your whole homework in one question- because i really am not trying to i just don't understand the question. Hope you answer the question and thank you in advance.
Answer by checkley77(12844)

40% * some amount of ml + 20% * (60-some amount of ml)=32% of 60 ml
.40x+.20(60-x)=.32*60
.40x+12-.20x=19.2
.20x=19.2-12
.20x=7.2
x=7.2/.20
x=36 ml of 40% solution is used.
60-36=24 ml of 20% solution is used.
Proof:
.40*36+.20*24=.32*60
14.4+4.8=19.2
19.2=19.2